The perimeter of a rectangle is 122 cm. The length is 11 cm less than five times the width. Find the dimensions of the rectangle.
Answer by TimothyLamb(4379)

p = 2L + 2w = 122
L = 5w - 11
---
2L + 2w = 122
2(5w - 11) + 2w = 122
10w - 22 + 2w = 122
12w = 144
---
w = 12
L = 49
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php